Q . FOR THE REACTION :-
   
   CaCO3 (s)------------ ------------> CaO (s) + CO2 (g),
                 Kp = 1.16 atm AT 8000C .
 
IF 20.0 g OF CaCO3  WERE KEPT IN A 10 L CONTAINER AND HEATED UP TO 8000 C , WHAT PERCENTAGE OF CaCO3 WOULD REMAIN UNREACTED AT EQUILIBRIUM.
                        
                                
                     
 
 
     

Grade:Select Grade

2 Answers

Naveen Kumar
9 years ago
CaCO3(s)------------------------> CaO (s) + CO2(g)
20g.....................................0g...........0g........(initially)
Kp=PCO2=1.16atm(as the solids are notincluded in the expression as their active mass is unity.)

Volume=10L, T=800K,So the number of moles of CO2 at equilibrium= can be calculated using ideal gas equation.
PV=nRT
n= {(1.16*10)/(0.082*800)} mol
and this much mole of CaCO3 would be consumeed .
So mass ofCaCO3=100g/mol *n
So mass left=20g-n*100g=l(let)
so % of CaCO3 unreacted=(l/20)*100
Bharat Makkar
9 years ago
thanks....